# Exponents and front fluctuations in the quenched Kardar-Parisi-Zhang universality class of one and two dimensional interfaces

### v1

@@ I Introduction
The collective properties that characterize dynamical complex systems often emerge from the interplay between external driving and dissipation, in such a way that frequently criticality holds spontaneously in large regions of parameter space [Grinstein1991, Odor2004, Pruessner12, Tauber14]. An example is surface kinetic roughening where, in principle, the physical object of interest is a rough surface or interface whose time evolution results from the interplay between deterministic and stochastic processes as for, e.g., thin film or epitaxial growth, or for bacterial colonies [Barabasi95, Krug97]. Notably, the critical behavior displayed by this type of systems is proving to be ubiquitous across system nature and physical scales, to the extent that some of its main instances —like the celebrated one-dimensional (1D) Kardar-Parisi-Zhang (KPZ) universality class [KPZ86, Kriecherbauer10, Halpin15, Takeuchi18]— are recently becoming relevant even to non-interfacial systems. Recent examples can be found from quantum [Fontaine2022, Wei2022, Sieberer2025] to active [Caballero2018b, Caballero2020] matter, the kinetics of chemical reactions [Mondal2022], or the synchronization of phase or limit cycle [Gutierrez2023, Gutierrez2023b, Gutierrez2024] oscillators.
@@
Actually, work on surface kinetic roughening, and in particular that related with the KPZ universality class [KPZ86, Halpin15, Takeuchi18], is recently expanding our understanding of critical phenomena at large. Indeed, the critical behavior that ensues is quite rich, as it implies universal behavior not only from the point of view of the values of the critical exponents, but also with respect to the probability distribution function (PDF) of one-point field fluctuations along the time-dependent regime of evolution [Kriecherbauer10, Halpin15, Takeuchi18]. Thus, for 1D KPZ interfaces, the PDF is a member of the celebrated Tracy-Widom (TW) family of probability distributions [Bertin2006, Makey20], which generalizes naturally into other PDF families in higher dimensions, as has been experimentally confirmed for turbulent liquid crystals [Takeuchi2010, Takeuchi18] and thin film growth [Almeida14, Halpin-Healy2014, Orrillo17] in the 1D and 2D cases, respectively. For the KPZ class, these PDF are conspicuously non-Gaussian, e.g., their non-zero skewness being related with the occurrence of a preferred growth direction [Krug97] and the irreversible character of the dynamics.
@@
Analogous behavior on the occurrence of universal PDF in the growth regime has been assessed for other universality classes as exemplified by the conserved KPZ equation [Carrasco2016], to the extent that, currently, the PDF of field fluctuations is considered as an important trait to unambiguously characterize the kinetic roughening universality class (see, for instance, recent discussions in Refs. [Rodriguez21, Marcos22] and others therein). Interestingly, this and other properties of the critical behavior of kinetically rough surfaces (like, for example, novel forms of dynamic scaling ansätze) have been uncovered for more standard dynamical critical systems only very recently, as in the non-conserved critical dynamics of the 2D Ising universality class [Vaquero2025], including experimental verification [Allemand2025].
@@
The kinetic roughening universality classes just mentioned correspond to systems in which the main source of randomness is time-dependent noise, whose competition with driving and dissipation at comparable time scales induces the so-called generic scale invariance alluded to above [Grinstein1991, Tauber14]. However, in many other contexts noise is quenched and does not evolve in the time scales of evolution for the interface, as in the physical examples of fluid flow or of an elastic manifold evolving in disordered media [Barabasi95]. In these systems, scale invariance also frequently ensues, with eventual connections to so-called self-organized criticality [Odor2004, Pruessner12, Wiese2022]. Back to surface kinetic roughening proper, one of the main universality classes in the presence of quenched disorder [Barabasi95, Odor2004, Wiese2022] is that of the quenched KPZ (qKPZ) equation [Galluccio1995, Stepanow1995],
@@
This model has a very rich dynamical behavior that has been fully elucidated recently in experiments of domain walls in magnetic films [Moon2013] and of reaction fronts in disordered media [Atis2015]. Here, we will concern ourselves with the $\lambda>0$ case only, [^1] For $\lambda<0$, Eq. (1) also describes a pinning transition, but for a very different class (so-called, negative qKPZ) of faceted interfaces [Jeong1996, Jeong1999, Atis2015]. where two different universality classes are identified. Thus, a pinning-depinning transition exists at a non-zero critical value of the driving force $F=F_{c}$, such that the interface is pinned, i.e., the average velocity is zero, or moving (non-zero average velocity) for $F\leq F_{c}$ or $F>F_{c}$, respectively [Barabasi95, Leschhorn1996, Ramasco01]. The front described by Eq. (1) only displays the exponent values that define the qKPZ universality class exactly at depinning when $F=F_{c}$ [Leschhorn1996]. For $F\gg F_{c}$ and $\lambda>0$, the scaling exponents are those of the (standard) KPZ universality class with time-dependent noise [Tang1995, Leschhorn1996]. For comparison, the scaling behavior of the latter KPZ equation [obtained by replacing the disorder $\eta(\boldsymbol{r},h)$ in Eq. (1) by similarly uncorrelated, Gaussian, time-dependent noise $\eta(\boldsymbol{r},t)$] depends neither on the value of $F$ nor on the sign of $\lambda$ [Barabasi95, Krug97].
Part of the interest generated by the qKPZ equation is due to its relation to directed percolation (DP), another paradigm of non-equilibrium phenomena [Henkel2008]. This connection was much developed early on through work on the discrete directed percolation depinning (DPD) model, believed to be in the qKPZ universality class and initially formulated to describe experiments on front propagation caused by fluid imbibition in paper [Amaral95, Barabasi95].
A mapping between the DPD model and DP allowed to interpret the critical exponents of DPD in terms of those of DP, in principle for any value of $d$. Much work has been devoted since to assessing the rich set of critical exponents and behaviors (like the statistics of avalanches [Chen2011], crossover behavior [Chen2015] and others) for this and other discrete models in the qKPZ universality class, and for the equation itself, see Ref. [Wiese2022] for a review. Very recently, further numerical and analytical works [Mukerjee2023, Mukerjee2023b] have revised previous arguments and exponent values in the literature [Amaral95], concluding in particular that the dynamic exponent $z$ does not derive directly from the assumed identification between the length of the invading path with the duration of an avalanche and reporting a different value for the exponent.
@@ II.1 The discrete model
The calculations reported herein have been performed on a lattice of length $L$ or on a square lattice of dimensions $L\times L$ for the $d=1$ and $d=2$ cases, respectively, subject to periodic boundary conditions (PBC). Following the discrete model (DM) put forward in Ref. [Song07], the local height $h(\boldsymbol{r},t)$ at a given time $t$ evolves according to
@@
where the sum spans the nearest neighbors of site $\boldsymbol{r}$, denoted as $\text{NN}(\boldsymbol{r})$. Finally, a centered-difference approach for the gradient has been taken, as in Refs. [Lee05, Song07, Mukerjee2023]. For the 1D model ($d=1$), we write
@@ II.2 Observables
where $F_{c}$ is the critical force, $\theta$ is the so-called velocity exponent, and we have neglected subdominant terms (scaling corrections). For $F=F_{c}$, on the other hand, the front velocity is a decreasing function of time, with the dominat term being given by [Lee05, Song07]
@@
where $\beta$ and $z$ are called growth and dynamic exponents, respectively. In Eq. (15), the scaling function $f(y)\sim\text{const.}$ for $t\ll L^{z}$, so that $w(t\ll L^{z})\sim t^{\beta}$, and $f(y)\sim y^{-\beta}$ for $t\gg L^{z}$, so that $w(t\gg L^{z})\sim\text{const.}\equiv w_{\mathrm{sat}}$. The saturation roughness scales with the size of the system, $L$, as $w_{\mathrm{sat}}\sim L^{\alpha}$, where $\alpha$ is the roughness exponent, and is reached after a saturation time $t_{\rm sat}\sim L^{z}$. Note that, as a consistency condition of the FV ansatz for the roughness, the $\alpha$, $\beta$, and $z$ exponents are not independent: Indeed, $L^{\alpha}\sim w_{\rm sat}\sim t_{\rm sat}^{\beta}\sim L^{z\beta}$, hence $\alpha=z\beta$ [Barabasi95].
Likewise, for the qKPZ universality class $\delta$ and $\beta$ are related to each other, since $\delta+\beta=1$ as suggested by an interpretation of $v(t)$ in terms of the density order parameter in phase transitions with absorbing states [Dickman2000], and verified for qKPZ at depinning [Lee05, Song07]. Indeed, provided that $\bar{h}(t)\sim w(t)$, then $1-\delta=\beta$, since $v\sim\bar{h}/t$ [Dickman2000, Barreales2023]. Note that this scaling relation is not satisfied, e.g., by the KPZ universality class with time-dependent noise, where $\delta=0$ but $\beta=1/3$, or by other kinetic roughening universality classes that feature non-trivial values of $\delta$. For instance, a continuum model for the spreading of a precursor film in band geometry features $\delta=0.42$ and $\beta=0.35$ [Marcos22].
As for the direct physical meaning of the kinetic roughening exponents, $\beta$ rules the growth with increasing time of the fluctuations of the front around the average front position, as evidenced by Eq. (14), while $\alpha$ is related to the fractal dimension of the front [Barabasi95, Mozo22]. Finally, the dynamic exponent $z$ is related to the time increase of the lateral correlation length $\xi(t)$ along the front, i.e.,
@@
The height-difference correlation function allows one to compute the correlation length $\xi(t)$ explicitly, as has been reported elsewhere [Barreales20]. Essentially, the correlation length can be thought of as the distance along the substrate coordinate at which the $C_{2}(\boldsymbol{r},t)$ function takes on a significant fraction of its value at the plateau value $C_{2}(L/2,t)$ (recall that we are considering PBC). Thus,
@@
where $a=0.9$ in this work; the scaling behavior of the correlation length is not much affected by the particular value of $a$ [Barreales20].
@@
Thus, fluctuations are centered around the average front and their time-dependent amplitude [as estimated by the roughness $w(t)$] is crucially normalized out. In this way, the ensuing PDF becomes stationary (i.e., the form of the PDF becomes time-independent) and universal in the sense that the same stationary PDF occurs for all members of the same kinetic roughening universality class, as has been verified in kinetic roughening systems [Kriecherbauer10, Halpin15, Takeuchi18, Carrasco2016, Barreales20, Rodriguez21, Marcos22, Barreales2023] and, recently, for more standard dynamical critical systems as well [Vaquero2025]. As noted above, to date the growth PDF is considered an important trait to unambiguously identify the kinetic roughening universality class, beyond the particular values of the scaling exponents.
The errors of all quantities reported in this work have been computed following the jackknife procedure, especially suitable for highly correlated data (e.g., all the fits performed in time) [Efron82, Young15]; see Appendix B of Ref. [Barreales20] for additional details. The system sizes, maximum times achieved, and number of realizations of the noise (or number of runs) considered for both models are collected in Table 5 of Appendix A.
@@ III.1 Evaluation of the critical force
The approach adopted in this work consists on evaluating the front velocity for various values of the external driving force $F$. According to Eq. (12), the fit of these results to a power law yields an estimation of both the critical force $F_{c}$ and the velocity exponent. Thus, Fig. 1 plots the velocity values vs. $F$, for different system sizes, computed for $d=1$ and $d=2$. In both cases, one observes that the front remains pinned (i.e., with zero velocity) for forces below some threshold which is dependent of the system size, specifically decreasing as $L$ increases.
For the 1D model [see Fig. 1(a)], we find $F_{c}=0.8111(3)$ and $\theta=0.673(6)$ for the largest size considered in this analysis $L=10^{6}$, [^2] For our largest value of the lattice size ($L_{\text{max}}=10^{6}$) we determine what values of the velocity are independent of the size of the system (i.e., $v$ is asymptotic, independent of $t$ and $L$): this happens for $F\geq F_{\text{min}}$, with $F_{\text{min}}$ depending on $L_{\text{max}}$. We perform fits to the scaling law, Eq. (12), in the interval $F_{\text{min}},F]$ starting from the higher simulated value of $F$ and reducing it until we obtain a $p$-value [Young15] larger than 5%. which we take as representative of the asymptotic limit. [^3] For uncorrelated data, all the fits reported in this paper have a $p$-value larger than 5%. Our values of $\theta$ and $F_{c}$ compare very well with those of Ref. [Song07], where $F_{c}=0.8105(5)$ and $\theta\approx 0.672$ were reported for the same model and disorder strength. Furthermore, our estimate of $\theta$ is consistent with the 1D DPD model, for which $\theta=0.66(4)$ [Amaral95]. Notice that, in principle, the value of $F_{c}$ is not expected to be universal, since it depends on the disorder strength implemented in the automaton model.
@@ III.2 Growth exponents
Table 1: Critical exponent values for the 1D model. We stress that the $\delta$, $\beta$, $z$, and $\alpha$ exponents have been computed directly (i.e., without using scaling relations). Besides these four exponents, we checked the scaling relations $\beta+\delta=1$ and $\beta z=\alpha$ (last two columns of the table). We recall that the value of $\alpha$ obtained using the scaling of the plateaus of the roughness is $\alpha=0.6325(17)$. We have been able to simulate, as a control, a small number of initial conditions (100) for $L=10^{7}$ in order to compute the $z$ and $\alpha$ exponents. The results obtained, marked with an asterisk in the table, are compatible with those for $L=10^{6}$, but with higher error for $z$. Therefore, our final estimate for $z$ will be that computed for $L=10^{6}$.
On the other hand, Fig. 3(a) displays the squared roughness data versus simulation time, as computed for different system sizes and $d=1$. In all cases a growth regime characterized by a well-defined value of $\beta$ follows an initial transient. The growth regime ends at maxima of the $w^{2}(t)$ curves, from where the squared width decreases to a time-independent saturation value. The fit of the numerical $w(t)$ data to Eq. (15), in the growth time interval where the condition $t\ll L^{z}$ is satisfied, yields the $\beta$ exponent values recorded in the third column of Table 1 for $d=1$. This exponent increases slightly with $L$, reaching a value $\beta=0.6290(6)$ for $L=10^{7}$. In relation to this, the sixth column of Table 1 shows that the sum $\beta+\delta$ is approaching one as $L$ increases. In particular, for $L=10^{7}$, we obtain $\beta+\delta=0.991(1)$ which is in good agreement with the theoretical expectation [Lee05, Song07].

### v2

@@ I Introduction
The collective properties that characterize dynamical complex systems often emerge from the interplay between external driving and dissipation, in such a way that frequently criticality holds spontaneously in large regions of parameter space [1, 2, 3, 4]. An example is surface kinetic roughening where, in principle, the physical object of interest is a rough surface or interface whose time evolution results from the interplay between deterministic and stochastic processes as for, e.g., thin film or epitaxial growth, or for bacterial colonies [5, 6]. Notably, the critical behavior displayed by this type of systems is proving to be ubiquitous across system nature and physical scales, to the extent that some of its main instances —like the celebrated one-dimensional (1D) Kardar-Parisi-Zhang (KPZ) universality class [7, 8, 9, 10]— are recently becoming relevant even to non-interfacial systems. Recent examples can be found from quantum [11, 12, 13] to active [14, 15] matter, the kinetics of chemical reactions [16], or the synchronization of phase or limit cycle [17, 18, 19] oscillators.
@@
Actually, work on surface kinetic roughening, and in particular that related with the KPZ universality class [7, 9, 10], is recently expanding our understanding of critical phenomena at large. Indeed, the critical behavior that ensues is quite rich, as it implies universal behavior not only from the point of view of the values of the critical exponents, but also with respect to the probability distribution function (PDF) of one-point field fluctuations along the time-dependent regime of evolution [8, 9, 10]. Thus, for 1D KPZ interfaces, the PDF is a member of the celebrated Tracy-Widom (TW) family of probability distributions [20, 21], which generalizes naturally into other PDF families in higher dimensions, as has been experimentally confirmed for turbulent liquid crystals [22, 10] and thin film growth [23, 24, 25] in the 1D and 2D cases, respectively. For the KPZ class, these PDF are conspicuously non-Gaussian, e.g., their non-zero skewness being related with the occurrence of a preferred growth direction [6] and the irreversible character of the dynamics.
@@
Analogous behavior on the occurrence of universal PDF in the growth regime has been assessed for other universality classes as exemplified by the conserved KPZ equation [26], to the extent that, currently, the PDF of field fluctuations is considered as an important trait to unambiguously characterize the kinetic roughening universality class (see, for instance, recent discussions in Refs. [27, 28] and others therein). Interestingly, this and other properties of the critical behavior of kinetically rough surfaces (like, for example, novel forms of dynamic scaling ansätze) have been uncovered for more standard dynamical critical systems only very recently, as in the non-conserved critical dynamics of the 2D Ising universality class [29], including experimental verification [30].
@@
The kinetic roughening universality classes just mentioned correspond to systems in which the main source of randomness is time-dependent noise, whose competition with driving and dissipation at comparable time scales induces the so-called generic scale invariance alluded to above [1, 4]. However, in many other contexts noise is quenched and does not evolve in the time scales of evolution for the interface, as in the physical examples of fluid flow or of an elastic manifold evolving in disordered media [5]. In these systems, scale invariance also frequently ensues, with eventual connections to so-called self-organized criticality [2, 3, 31]. Back to surface kinetic roughening proper, one of the main universality classes in the presence of quenched disorder [5, 2, 31] is that of the quenched KPZ (qKPZ) equation [32, 33],
@@
This model has a very rich dynamical behavior that has been fully elucidated recently in experiments of domain walls in magnetic films [34] and of reaction fronts in disordered media [35]. Here, we will concern ourselves with the $\lambda>0$ case only, [^1] For $\lambda<0$, Eq. (1) also describes a pinning transition, but for a very different class (so-called, negative qKPZ) of faceted interfaces [36, 37, 35]. where two different universality classes are identified. Thus, a pinning-depinning transition exists at a non-zero critical value of the driving force $F=F_{c}$, such that the interface is pinned, i.e., the average velocity is zero, or moving (non-zero average velocity) for $F\leq F_{c}$ or $F>F_{c}$, respectively [5, 38, 39]. The front described by Eq. (1) only displays the exponent values that define the qKPZ universality class exactly at depinning when $F=F_{c}$ [38]. For $F\gg F_{c}$ and $\lambda>0$, the scaling exponents are those of the (standard) KPZ universality class with time-dependent noise [40, 38]. For comparison, the scaling behavior of the latter KPZ equation [obtained by replacing the disorder $\eta(\boldsymbol{r},h)$ in Eq. (1) by similarly uncorrelated, Gaussian, time-dependent noise $\eta(\boldsymbol{r},t)$] depends neither on the value of $F$ nor on the sign of $\lambda$ [5, 6].
Part of the interest generated by the qKPZ equation is due to its relation to directed percolation (DP), another paradigm of non-equilibrium phenomena [41]. This connection was much developed early on through work on the discrete directed percolation depinning (DPD) model, believed to be in the qKPZ universality class and initially formulated to describe experiments on front propagation caused by fluid imbibition in paper [42, 5].
A mapping between the DPD model and DP allowed to interpret the critical exponents of DPD in terms of those of DP, in principle for any value of $d$. Much work has been devoted since to assessing the rich set of critical exponents and behaviors (like the statistics of avalanches [43], crossover behavior [44] and others) for this and other discrete models in the qKPZ universality class, and for the equation itself, see Ref. [31] for a review. Very recently, further numerical and analytical works [45, 46] have revised previous arguments and exponent values in the literature [42], concluding in particular that the dynamic exponent $z$ does not derive directly from the assumed identification between the length of the invading path with the duration of an avalanche and reporting a different value for the exponent.
@@ II.1 The discrete model
The calculations reported herein have been performed on a lattice of length $L$ or on a square lattice of dimensions $L\times L$ for the $d=1$ and $d=2$ cases, respectively, subject to periodic boundary conditions (PBC). Following the discrete model (DM) put forward in Ref. [47], the local height $h(\boldsymbol{r},t)$ at a given time $t$ evolves according to
@@
where the sum spans the nearest neighbors of site $\boldsymbol{r}$, denoted as $\text{NN}(\boldsymbol{r})$. Finally, a centered-difference approach for the gradient has been taken, as in Refs. [48, 47, 45]. For the 1D model ($d=1$), we write
@@ II.2 Observables
where $F_{c}$ is the critical force, $\theta$ is the so-called velocity exponent, and we have neglected subdominant terms (scaling corrections). For $F=F_{c}$, on the other hand, the front velocity is a decreasing function of time, with the dominant term being given by [48, 47]
@@
where $\beta$ and $z$ are called growth and dynamic exponents, respectively. In Eq. (15), the scaling function $f(y)\sim\text{const.}$ for $t\ll L^{z}$, so that $w(t\ll L^{z})\sim t^{\beta}$, and $f(y)\sim y^{-\beta}$ for $t\gg L^{z}$, so that $w(t\gg L^{z})\sim\text{const.}\equiv w_{\mathrm{sat}}$. The saturation roughness scales with the size of the system, $L$, as $w_{\mathrm{sat}}\sim L^{\alpha}$, where $\alpha$ is the roughness exponent, and is reached after a saturation time $t_{\rm sat}\sim L^{z}$. Note that, as a consistency condition of the FV ansatz for the roughness, the $\alpha$, $\beta$, and $z$ exponents are not independent: Indeed, $L^{\alpha}\sim w_{\rm sat}\sim t_{\rm sat}^{\beta}\sim L^{z\beta}$, hence $\alpha=z\beta$ [5].
Likewise, for the qKPZ universality class $\delta$ and $\beta$ are related to each other, since $\delta+\beta=1$ as suggested by an interpretation of $v(t)$ in terms of the density order parameter in phase transitions with absorbing states [49], and verified for qKPZ at depinning [48, 47]. Indeed, provided that $\bar{h}(t)\sim w(t)$, then $1-\delta=\beta$, since $v\sim\bar{h}/t$ [49, 50]. Note that this scaling relation is not satisfied, e.g., by the KPZ universality class with time-dependent noise, where $\delta=0$ but $\beta=1/3$, or by other kinetic roughening universality classes that feature non-trivial values of $\delta$. For instance, a continuum model for the spreading of a precursor film in band geometry features $\delta=0.42$ and $\beta=0.35$ [28].
As for the direct physical meaning of the kinetic roughening exponents, $\beta$ rules the growth with increasing time of the fluctuations of the front around the average front position, as evidenced by Eq. (14), while $\alpha$ is related to the fractal dimension of the front [5, 51]. Finally, the dynamic exponent $z$ is related to the time increase of the lateral correlation length $\xi(t)$ along the front, i.e.,
@@
The height-difference correlation function allows one to compute the correlation length $\xi(t)$ explicitly, as has been reported elsewhere [52]. Essentially, the correlation length can be thought of as the distance along the substrate coordinate at which the $C_{2}(\boldsymbol{r},t)$ function takes on a significant fraction of its value at the plateau value $C_{2}(L/2,t)$ (recall that we are considering PBC). Thus,
@@
where $a=0.9$ in this work; the scaling behavior of the correlation length is not much affected by the particular value of $a$ [52].
@@
Thus, fluctuations are centered around the average front and their time-dependent amplitude [as estimated by the roughness $w(t)$] is crucially normalized out. In this way, the ensuing PDF becomes stationary (i.e., the form of the PDF becomes time-independent) and universal in the sense that the same stationary PDF occurs for all members of the same kinetic roughening universality class, as has been verified in kinetic roughening systems [8, 9, 10, 26, 52, 27, 28, 50] and, recently, for more standard dynamical critical systems as well [29]. As noted above, to date the growth PDF is considered an important trait to unambiguously identify the kinetic roughening universality class, beyond the particular values of the scaling exponents.
The errors of all quantities reported in this work have been computed following the jackknife procedure, especially suitable for highly correlated data (e.g., all the fits performed in time) [53, 54]; see Refs [55, 56] and Appendix B of Ref. [52] for additional details. The system sizes, maximum times achieved, and number of realizations of the noise (or number of runs) considered for both models are collected in Table 5 of Appendix A.
@@ III.1 Evaluation of the critical force
The approach adopted in this work consists of evaluating the front velocity for various values of the external driving force $F$. According to Eq. (12), the fit of these results to a power law yields an estimation of both the critical force $F_{c}$ and the velocity exponent. Thus, Fig. 1 plots the velocity values vs. $F$, for different system sizes, computed for $d=1$ and $d=2$. In both cases, one observes that the front remains pinned (i.e., with zero velocity) for forces below some threshold which is dependent on the system size, specifically decreasing as $L$ increases.
For the 1D model [see Fig. 1(a)], we find $F_{c}=0.8111(3)$ and $\theta=0.673(6)$ for the largest size considered in this analysis $L=10^{6}$,  [^2] For our largest value of the lattice size ($L_{\text{max}}=10^{6}$) we determine what values of the velocity are independent of the size of the system (i.e., $v$ is asymptotic, independent of $t$ and $L$): this happens for $F\geq F_{\text{min}}$, with $F_{\text{min}}$ depending on $L_{\text{max}}$. We perform fits to the scaling law, Eq. (12), in the interval $F_{\text{min}},F]$ starting from the higher simulated value of $F$ and reducing it until we obtain a $p$-value [54] larger than 5%. We then continue the analysis by further increasing the minimum force in order to verify the stability of the exponent $\theta$ and the critical force $F_{c}$. which we take as representative of the asymptotic limit. [^3] For uncorrelated data, all the fits reported in this paper have a $p$-value larger than 5%. Our values of $\theta$ and $F_{c}$ compare very well with those of Ref. [47], where $F_{c}=0.8105(5)$ and $\theta\approx 0.672$ were reported for the same model and disorder strength. Furthermore, our estimate of $\theta$ is consistent with the 1D DPD model, for which $\theta=0.66(4)$ [42]. Notice that, in principle, the value of $F_{c}$ is not expected to be universal, since it depends on the disorder strength implemented in the automaton model.
@@ III.2 Growth exponents
Table 1: Critical exponent values for the 1D model. We stress that the $\delta$, $\beta$, $z$, and $\alpha$ exponents have been computed directly (i.e., without using scaling relations). Besides these four exponents, we checked the scaling relations $\beta+\delta=1$ and $\beta z=\alpha$ (last two columns of the table). We recall that the value of $\alpha$ obtained using the scaling of the plateaus of the roughness is $\alpha=0.6325(17)$. We have been able to simulate, as a control, a small number of initial conditions (100) for $L=10^{7}$ in order to compute the $z$ and $\alpha$ exponents. The results marked with an asterisk are compatible with those for $L=10^{6}$, but have a higher error for $z$. Therefore, our final estimate for $z$ is that computed for $L=10^{6}$.
On the other hand, Fig. 3(a) displays the squared roughness data versus simulation time, as computed for different system sizes and $d=1$. In all cases a growth regime characterized by a well-defined value of $\beta$ follows an initial transient. The growth regime ends at maxima of the $w^{2}(t)$ curves, from where the squared width decreases to a time-independent saturation value. The fit of the numerical $w(t)$ data to Eq. (15), in the growth time interval where the condition $t\ll L^{z}$ is satisfied, yields the $\beta$ exponent values recorded in the third column of Table 1 for $d=1$. This exponent increases slightly with $L$, reaching a value $\beta=0.6290(6)$ for $L=10^{7}$. In relation to this, the sixth column of Table 1 shows that the sum $\beta+\delta$ is approaching one as $L$ increases. In particular, for $L=10^{7}$, we obtain $\beta+\delta=0.991(1)$ which is in good agreement with the theoretical expectation [48, 47].